OpenStudy (anonymous):

In a mathematics class, the ratio of women to men is 6 to 5. If there are 44 students in the class, how many men are there?

OpenStudy (tkhunny):

You don't ALWAYS need a fancy formula. Sometimes, you can just count. 6 + 5 = 11 12 + 10 = 22 24 + 20 = 44

OpenStudy (anonymous):

Ok, how do you get the 12 + 10 and the 24 + 20? @tkhunny

OpenStudy (tkhunny):

Just double everything. See how the ratio is unaffected? \(\dfrac{24}{20} = \dfrac{12}{10} = \dfrac{6}{5}\)